# Reinforced Imitative Trajectory Planning for Urban Automated Driving

### v1

@@ 5.5 Qualitative Results
Some qualitative results on the testing set are presented in this section. Fig. 8 depicts a scenario of traversing an intersection, where the expert turned right with a trajectory represented by the orange curves in Fig. 8(a), (b), and (c). The ego vehicle controlled by UrbanDriver, QCMAE, and HRITP produced different trajectories of the ego vehicle as depicted by the blue curves in Fig. 8(a), (b), and (c). The top and bottom halves of Fig. 8(a), (b), and (c) represent the first and last frames of simulation, respectively. The motion states are shown in Fig. 8(d). As shown in Fig. 8, at the beginning of the simulation, UrbanDriver and QCMAE were able to drive within the lane. However, they drove out of the drivable area later, which was caused by covariate shift. In contrast, the ego vehicle controlled by HRITP entered the intersection following the preceding vehicle and properly stopped when the preceding vehicle stopped, probably due to pick-up or drop-off, just as the expert did.
@@
Another scenario is shown in Fig. 9, where the ego vehicle was supposed to stop near the intersection. However, the UrbanDriver quickly veered off, running a red light and even driving into the wrong lane, as depicted in Fig. 9(a). QCMAE performed much better during the first 7 seconds but gradually drove out of the drivable area over the next 8 seconds, as presented in Fig. 9(b). By comparison, HRITP successfully guided the ego vehicle to slow down and keep a safe distance from the preceding vehicle, as shown in Fig. 9(c).
@@ 5.6 Discussion
The application of RL in trajectory planning for urban automated driving is limited by the poor convergence of RL and the difficulty of designing reward functions. The proposed RITP method integrates RL with IL to overcome the convergence problem. It trains CriticFormer to evaluate the closed-loop effect of an action, according to which supervisory signals are generated for IL. In contrast to RL, the gradients computed by IL are sufficient to optimize deep neural networks such as MotionFormer. Then, MotionFormer trained through IL explores the state and action spaces using the trajectory noise, thus resolving the suboptimality and the online computational complexity issues in TRAVL. Moreover, Bayesian RewardFormer provides effective reward signals for RL. As indicated by the results presented in the previous subsections, RITP significantly outperforms the baselines in terms of the closed-loop metrics. Another important reason for the effectiveness of RITP is that CriticFormer, MotionFormer, and Bayesian RewardFormer utilize the attention mechanism to extract information from the query-centric representation effectively. We also found that combining data-driven and model-driven methods can further improve the performance of RITP. Due to the inevitable epistemic uncertainty in neural networks, RITP might plan dangerous trajectories in unfamiliar situations. The post-optimizer refines the trajectories output by RITP to correct these wrong plans.
It is worth noting that common IL-based methods can be incorporated into the proposed RITP framework by simply replacing MotionFormer and its training objective. For example, MotionFormer can be replaced by UniAD model [3], which takes as input multi-view camera images, to enable end-to-end trajectory planning [13]. With RL, the generalization ability of these models can be enhanced to accomplish safe trajectory planning in urban scenarios.
@@ 6 Conclusion
This paper proposes a novel RL-based trajectory planning method for urban automated driving. The proposed method integrates IL into RL to improve the convergence, thus enabling planning for multiple future steps. Furthermore, a transformer-based Bayesian reward function is developed, which removes the linear structure assumption imposed by most existing methods and is applicable to RL in urban scenarios. Besides, a hybrid-driven trajectory planning framework is proposed to enhance the feasibility of the trajectories planned by the RL agent. The results of the experiments conducted on the large-scale real-world urban automated driving nuPlan dataset indicate that the proposed methods significantly outperform baselines in terms of the closed-loop metrics.
Nomenclature
@@
HRITP Hybrid-driven reinforced
imitative trajectory planning

### v2

@@ 5.5 Qualitative Results
This section presents some qualitative results on the testing set. Fig. 8 illustrates a scenario involving traversing an intersection. Fig. 8(a)-(d) depict the expert’s right turn trajectory (orange curves) and the corresponding ego vehicle trajectories generated by UrbanDriver, QCMAE, RITP, and PLUTO (blue curves). In these subfigures, the top and bottom halves represent the first and last frames of the simulation, respectively. The motion states (e.g., speed, acceleration) of the ego vehicle for each method are presented in Fig. 8(e). As shown in Fig. 8, UrbanDriver and QCMAE were able to drive within the lane at the beginning of the simulation. However, they subsequently departed from the drivable area, which was caused by covariate shift. In contrast, the ego vehicle controlled by RITP followed the preceding vehicle into the intersection and stopped appropriately when the preceding vehicle halted (possibly for pick-up or drop-off), effectively mimicking the expert’s behavior. PLUTO also succeeded in navigating this scenario, although it exhibited less stable acceleration, as evident in Fig. 8(e).
@@
Fig. 9 illustrates another scenario, where the ego vehicle was supposed to stop near the intersection. However, as depicted in Fig. 9(a), the UrbanDriver quickly veered off, running a red light and even driving into the wrong lane. QCMAE performed significantly better during the initial 7 seconds but subsequently drove out of the drivable area over the following 8 seconds, as shown in Fig. 9(b). In contrast, RITP successfully guided the ego vehicle to slow down and maintain a safe distance from the preceding vehicle, while PLUTO planned reasonable lane change maneuvers, as shown in Fig. 9(c) and (d).
@@
5.6 Ablation Study
Closed-loop simulations were conducted on the validation set, covering both non-reactive and reactive modes, to analyze the impact of the following factors on model performance: the multi-step reasoning (MSR), the warm start (WS), the multi-agent supervision (MAS, see Section 2.8), and the hybrid-driven planning framework (Hybrid, see Section 4). Specifically, MSR refers to training MotionFormer to plan an 8-second trajectory (80 future positions), whereas without MSR, it was trained for a 1-second trajectory (2 future positions). WS indicates that MotionFormer was initialized with weights from the trained QCMAE model, whereas without WS, it was trained from scratch. Note that the standard TD3 [27] algorithm was employed to train MotionFormer without MSR, while our RITP method was used in the other ablation experiments.
Table 3 summarizes the results of the ablation study. As shown in Table 3, $\mathcal{M}_{0}$, which omits MSR, WS, MAS, and Hybrid planning, exhibits the poorest performance. Although $\mathcal{M}_{1}$ incorporated MSR, it also has poor closed-loop planning performance. The performance of $\mathcal{M}_{2}$ highlights the necessity of warm-starting the training with the trained QCMAE model. Furthermore, incorporating MAS ($\mathcal{M}_{3}$) results in a significant improvement in closed-loop non-reactive and reactive scores (39.4% and 45.7%, respectively) compared to approaches that solely optimize the ego-vehicle trajectory outputted by MotionFormer. This significant improvement is attributed to MAS simultaneously optimizing the predicted trajectories of all surrounding vehicles, thereby increasing data richness and facilitating better trajectory modeling. Moreover, with the hybrid planning framework, $\mathcal{M}_{4}$ further significantly enhances planning performance across most metrics compared to $\mathcal{M}_{3}$, with an acceptable reduction observed in CMFT. Specifically, the closed-loop non-reactive and reactive scores of $\mathcal{M}_{4}$ improved by 82.6% and 62.7% compared to $\mathcal{M}_{3}$.
Table 3: Ablation study of the impact of each factor on model performance on the validation set
Mode Model MSR WS MAS Hybrid NAFC TTCWB DAC DDC SLC MP PARR CMFT SR score
CL-NR $\mathcal{M}_{0}$ 0.30 0.26 0.06 0.95 0.99 0.47 0.40 0.60 0.31 0.02
$\mathcal{M}_{1}$checkmark 0.31 0.29 0.27 0.89 0.93 0.88 0.74 0.70 0.32 0.13
$\mathcal{M}_{2}$checkmark $\checkmark$ 0.70 0.62 0.79 0.99 0.99 0.61 0.46 0.80 0.71 0.33
$\mathcal{M}_{3}$checkmark $\checkmark$ $\checkmark$ 0.76 0.72 0.78 0.99 0.99 0.78 0.59 0.97 0.77 0.46
$\mathcal{M}_{4}$checkmark $\checkmark$ $\checkmark$ $\checkmark$ 0.96 0.91 0.97 1.00 1.00 0.96 0.85 0.84 0.96 0.84
CL-R $\mathcal{M}_{0}$ 0.37 0.33 0.08 0.95 0.99 0.47 0.40 0.61 0.38 0.04
$\mathcal{M}_{1}$checkmark 0.38 0.32 0.30 0.88 0.93 0.88 0.74 0.70 0.39 0.14
$\mathcal{M}_{2}$checkmark $\checkmark$ 0.76 0.71 0.78 0.99 0.99 0.60 0.46 0.78 0.77 0.35
$\mathcal{M}_{3}$checkmark $\checkmark$ $\checkmark$ 0.84 0.80 0.80 0.99 0.99 0.77 0.57 0.98 0.84 0.51
$\mathcal{M}_{4}$checkmark $\checkmark$ $\checkmark$ $\checkmark$ 0.96 0.92 0.98 1.00 1.00 0.97 0.81 0.84 0.97 0.83
We also attempted to train $\mathcal{M}_{1}$and $\mathcal{M}_{2}$ using the standard TD3 algorithm instead of our RITP method. However, the training failed to converge, and the resulting model yielded erratic trajectories consisting of a set of disorganized positions.
5.7 Runtime Analysis
This section provides an analysis of the computational time complexity and runtime of the proposed method. During inference, the computational time complexity of RITP stems primarily from the attention modules within MotionFormer and the post-optimizer components: trajectory proposal generation, QP path planning, and QP speed planning. Given the parameters $A$ (number of agents), $T$ (number of historical time steps), and $M$ (number of map polygons), the time complexity of the attention mechanisms dominates, specifically: temporal attention ($O(AT^{2})$), map-map attention ($O(M^{2})$), agent-map attention ($O(ATM)$), and agent-agent attention ($O(A^{2}T)$). The post-optimizer involves curve interpolation, LQR, and QP. These sub-problems can be solved efficiently using mature mathematical tools and have been proven to meet the real-time requirements of automated vehicles, as demonstrated by their use in systems like Baidu Apollo [50].
Table 4 presents the model inference time (mean plus/minus one standard deviation) in a dense traffic scenario involving 71 agents and 240 map polygons. The model was run on a single NVIDIA RTX 4090 GPU and an Intel Xeon 4310 processor (up to 3.3 GHz). The results indicate that the MotionFormer component exhibits very low latency. This efficiency is primarily attributable to the highly parallel CUDA architecture and dedicated Tensor Cores of the NVIDIA GPU, along with its efficient support for low-precision inference and batch computation. In contrast, the Post-Optimizer component incurs higher runtime. This is because the Post-Optimizer was implemented in Python, an interpreted language known for significantly lower execution efficiency compared to C++. Based on this, it is reasonable to expect that reimplementing the Post-Optimizer in C++ would yield significant computational efficiency improvements.
Table 4: Model inference time on the validation set
Component Online Inference time (ms)
MotionFormer 56 ($\pm$ 7)
Post-Optimizer 2616 ($\pm$ 292)
Total 2672 ($\pm$ 297)
@@ 5.8 Discussion
The application of RL in trajectory planning for urban automated driving is limited by the poor convergence of RL and the difficulty of designing reward functions. The proposed RITP method integrates RL with IL to overcome the convergence problem. It trains CriticFormer to evaluate the closed-loop effect of an action, according to which supervisory signals are generated for IL. In contrast to RL, the gradients computed by IL are sufficient to optimize deep neural networks such as MotionFormer. Then, MotionFormer trained through IL explores the state and action spaces using the trajectory noise, thus resolving the suboptimality and the online computational complexity issues in TRAVL. Moreover, Bayesian RewardFormer provides effective reward signals for RL. As indicated by the results presented in the previous subsections, RITP significantly outperforms the baselines in terms of the closed-loop metrics. Another important reason for the effectiveness of RITP is that CriticFormer, MotionFormer, and Bayesian RewardFormer utilize the attention mechanism to extract information from the query-centric representation effectively. We also found that combining data-driven and model-driven methods can further improve the performance of RITP. Due to the inevitable epistemic uncertainty in neural networks, RITP might plan dangerous trajectories in unfamiliar situations. The post-optimizer refines the trajectories outputted by RITP to correct these wrong plans.
It is worth noting that common IL-based methods can be incorporated into the proposed RITP framework by simply replacing MotionFormer module and its associated training objective. For example, MotionFormer can be replaced by UniAD model [3], which takes as input multi-view camera images, to enable end-to-end trajectory planning [13]. By leveraging the proposed RL framework, the generalization ability of such incorporated models can be significantly enhanced for safe trajectory planning in complex urban scenarios. Beyond model replacement, the RITP framework’s training objective could potentially be enhanced by integrating SOTA IL techniques. For instance, incorporating contrastive imitation learning, as employed in PLUTO [42], holds potential for further performance improvement.
@@ 6 Conclusion
This paper proposes a novel RL-based trajectory planning method for urban automated driving. The proposed method integrates IL into RL to improve the convergence, thus enabling planning for multiple future steps. Furthermore, a transformer-based Bayesian reward function is developed, which removes the linear structure assumption imposed by most existing methods and is applicable to RL in urban scenarios. Besides, a hybrid-driven trajectory planning framework is proposed to enhance the feasibility of the trajectories planned by the RL agent. The results of the experiments conducted on the large-scale real-world urban automated driving nuPlan dataset indicate that the proposed method significantly outperforms baselines with identical policy network structures and achieves competitive performance compared to the SOTA method in terms of closed-loop metrics.
Abbreviations